# Probing of EoS with clusters and hypernuclei

### v1

@@ VI.1 Mean Transverse Momentum
Figure 15: The mean transverse momentum $\langle p_{T}\rangle$ of protons and $\Lambda$s as a function of rapidity (left panels), and at mid-rapidity as a function of the number of participating nucleons (right panels) in $\sqrt{s_{\rm{NN}}}=3$ Au$+$Au GeV collisions [11, 13]. The data are shown as solid markers, while PHQMD calculations are shown as lines.
Fig. 15 shows $\langle p_{T}\rangle$ of $p$ and $\Lambda$ as a function of rapidity on the left and for the rapidity interval $-0.1<y<0$ as a function of the number of participants on the right hand side. The number of participants for each centrality is determined by the same Glauber calculations, which we used for the centrality determination, described in Sec. IV. Calculations with a S EoS underestimate the data for both, $p$ and $\Lambda$. As a function of rapidity, calculations using a S EoS reproduce the same functional form as the data for both $p$ and $\Lambda$. In centrality dependence, however, the discrepancy between data and calculations increases with centrality for $p$, whereas for $\Lambda$, the calculations follow the experimental trend which flattens with increasing centrality.
@@
It is worthwhile to investigate this further. To separate the influence from the temperature and the collective expansion on the transverse momentum spectra, we perform blast-wave fits to the proton and $\Lambda$ spectra in the rapidity region $y=(-0.1,0)$ in four different centralities, $0-10\%,10-20\%,20-40\%$, and $40-80\%$ to extract the temperature at kinetic freeze out $T_{kin}$ and the average collective velocity $\langle\beta_{T}\rangle$. The radial flow profile parameter $n$ is fixed to 1 in these studies. The results are shown in Fig. 16. The temperature of protons and $\Lambda$s are identical within uncertainties but the collective radial flow of protons is systematically larger than that of $\Lambda$s. $\Lambda$s are produced particles and the production probability is higher close to the center of the fireball where the nucleon density is high and consequently the mean free path small. In the center of the fireball the collective velocity is small (see eq. 30) which results in a lower collective velocity of the $\Lambda$s compared to protons. The blast-wave fit results over the full $p_{T}$ range from PHQMD SM simulations are shown in the same figure as colored bands. When performing the blast-wave fits to the PHQMD $p$ and $\Lambda$ $p_{T}$ spectra, an additional 5% uncertainty (fudge factor) was added to the statistical errors. This adjustment was applied to achieve a reasonable $\chi^{2}/\mathrm{ndf}$ and to match the level of uncertainty of the experimental data [11, 13]. PHQMD describes the trends observed in the data well, and the deviation seen in peripheral collisions for protons has been discussed in the previous section.
@@ VI.2 Directed Flow
Figure 20, bottom panel, shows that the values of $v_{1}/A$ of all light nuclei, including protons, approximately follows an A scaling for $-0.3<y<0$. This is is also seen in PHQMD calculations. For all cluster sizes the theoretical $v_{1}(y)$ and $v_{1}(p_{T})$ calculations with a H EoS provide the best description of the data. SM slightly underpredicts the data, while a S EOS substantially deviates from the experimental data.
In Fig. 21 we present the PHQMD results for $v_{1}(y)$ of protons (top left), $\Lambda$ (top right) in comparison to the STAR data [14] for $\sqrt{s_{\rm{NN}}}=3$ GeV $10-40\%$ central Au+Au collisions and for the $p_{T}$ interval $2.00.4<p_{T}<$ 2.0 GeV$/c$ for a hard, soft and momentum-dependent EoS. In the bottom figure we display $v_{1}(y)$ of protons and $\Lambda$s together. $v_{1}(y)$ of protons and $\Lambda$s is negative. It is remarkable that the data for both $\Lambda$ and proton are similar despite of the quite different radial expansion velocities of the blast wave fit.
We also note that although $v_{1}$ of protons in H and SM calculations are similar, it differs quite significantly for $\Lambda$. In PHQMD the results differ because the protons close to the transverse surface of the fireball obtain more directed flow than those close to the center of the fireball [71]. Since $\Lambda$s have a higher chance to be produced in the center of the fireball than close to the surface, their flow (as well as that of $K^{+}$, the produced associated meson), is smaller. This may indicate that the parametrization of the SM EoS may be too simplistic in being not stiff enough at high densities or that the $\Lambda N$ scattering is underestimated, although in PHQMD the experimentally measured (free) $\Lambda N$ cross section is implemented. A detailed investigation of the directed flow of all strange hadrons may elucidate this interesting issue. We observe that calculations with a S EoS, similar to the comparisons of the $\langle p_{T}\rangle$, do not provide a good description of the data. They underestimate the $v_{1}$ for both $\Lambda$ and proton in all measured rapidity ranges. However, although calculations with SM, which provides good description of $\langle p_{T}\rangle$, describe the $v_{1}$ of $p$ fairly well, they overestimate the $v_{1}$ of $\Lambda$. The flow is also influenced by the choice of the $\Lambda$ potential [72], which we assume here to be identical to the potential of nucleons.
Figure 19: The directed flow $v_{1}$ of proton, deuteron, triton, ${}^{3}\rm{He}$, and ${}^{4}\rm{He}$ as a function of rapidity in $10-40\%$ mid-central Au$+$Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. The STAR data [14] are shown as colored markers while PHQMD calculations are shown as colored lines or bands.
Figure 20: The directed flow $v_{1}$ of proton, deuteron, triton, ${}^{3}\rm{He}$, and ${}^{4}\rm{He}$ as a function of $p_{T}$ in different rapidity intervals in $10-40\%$ mid-central Au$+$Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. The STAR data [12] are shown as black markers while PHQMD calculations are shown as colored bands.
Figure 21: The directed flow $v_{1}$ of proton and $\Lambda$ as a function of rapidity in $10-40\%$ mid-central $\sqrt{s_{\rm{NN}}}=3$ GeV Au$+$Au collisions. The STAR data [12] are shown as black closed (proton) and open ($\Lambda$) markers while PHQMD calculations are shown as colored bands.
@@ VI.3 Elliptic Flow
Now we step to a comparison of the PHQMD results for the elliptic flow for Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV with the STAR data [12]. In Fig. 22 we show the elliptic flow ($v_{2}$) of protons, deuterons and tritons, ${}^{3}\rm{He}$, and ${}^{4}\rm{He}$ as a function of the rapidity $y$ and integrated in the chosen $p_{T}$ ranges for the three EoS. For protons, $v_{2}$ is slightly negative in the considered $p_{T}$ interval. One can see that the $v_{2}$ of protons is quite well described by the SM EoS. The H EoS predicts almost zero $v_{2}$, while the S EoS predicts a positive $v_{2}$, which is opposite to the data. For deuterons, tritons, ${}^{3}\rm{He}$, and ${}^{4}\rm{He}$, the soft EoS yields a positive $v_{2}(y)$, in contradistinction to the negative $v_{2}$ of the experimental data. For all clusters SM and H EOS give rather similar results . Both show a more negative $v_{2}$ than the data in the whole rapidity range in distinction to the PHQMD calculations at $\sqrt{s_{\rm{NN}}}=2.4$ GeV [31], where the $v_{2}$ for SM and H is less negative than the data, however in a slightly different $p_{T}$ and centrality interval. It is remarkable that the sign of $v_{2}$ depends on the cluster size. For protons, $v_{2}$ is always negative, for $d$, $t$ and ${}^{3}{\rm{He}}$ it changes from positive values at $y=-0.5$ to negative values close to midrapidity, whereas for heavier fragments $v_{2}$ is always positive. PHQMD calculations with H and SM reproduce the trend of $v_{2}$ as a function of the rapidity.
@@
In the bottom right plot of Fig. 22 we compile $v_{2}(y)$ of the experimental results and the theoretical predictions with a SM EoS for protons, deuterons, triton, ${}^{3}\rm{He}$, and ${}^{4}\rm{He}$. The rapidity distributions of $v_{2}$ for different clusters exhibit a clear mass ordering, similar to $v_{1}$. This is qualitatively reproduced by the PHQMD calculations using a H or SM EoS. At mid-rapidity, $-0.1<y<0$, the value of $v_{2}$ is negative and nearly identical for $p$, $d$, and ${}^{3}\rm{He}$. A soft EoS give, as said, positive $v_{2}$ values for all clusters in $-0.5<y<0$, opposite to the data.

### v2

@@ VI.1 Mean Transverse Momentum
$\langle p_{T}\rangle(T,\beta)=\frac{\int p_{T}\frac{dN}{dp_{T}}dp_{T}}{\int\frac{dN}{dp_{T}}dp_{T}}.$ (32)
@@
Fig. 14 shows $\langle p_{T}\rangle$ of $p$ and $\Lambda$ as a function of rapidity on the left and for the rapidity interval 0.1<y<0-0.1<y<0 as a function of the number of participants on the right hand side. The number of participants for each centrality is determined by the same Glauber calculations, which we used for the centrality determination, described in Sec. IV. Calculations with a S EoS underestimate the data for both, $p$ and $\Lambda$. As a function of rapidity, calculations using a S EoS reproduce the same functional form as the data for both $p$ and $\Lambda$. In centrality dependence, however, the discrepancy between data and calculations increases with centrality for $p$, whereas for $\Lambda$, the calculations follow the experimental trend which flattens with increasing centrality.
@@
[…]
@@ VI.2 Directed Flow
Figure 18: The directed flow $v_{1}$ of proton, deuteron, triton, $^{3}\rm{He}$, and $^{4}\rm{He}$ as a function of rapidity in $10-40\%$ mid-central Au++Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. The STAR data [14] are shown as colored markers while PHQMD calculations are shown as colored lines or bands.
@@
Figure 19: The directed flow $v_{1}$ of proton, deuteron, triton, $^{3}\rm{He}$, and $^{4}\rm{He}$ as a function of $p_{T}$ in different rapidity intervals in $10-40\%$ mid-central Au++Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. The STAR data [12] are shown as black markers while PHQMD calculations are shown as colored bands.
Figure 19, bottom right panel, shows that the values of $v_{1}/A$ of all light nuclei, including protons, approximately follows an A scaling for 0.3<y<0-0.3<y<0. This is is also seen in PHQMD calculations. For all cluster sizes the theoretical $v_{1}(y)$ and $v_{1}(p_{T})$ calculations with a H EoS provide the best description of the data. SM slightly underpredicts the data, while a S EOS substantially deviates from the experimental data.
[…]
Figure 20: The directed flow $v_{1}$ of proton and $\Lambda$ as a function of rapidity in $10-40\%$ mid-central $\sqrt{s_{\rm{NN}}}=3$ GeV Au++Au collisions. The STAR data [12] are shown as black markers while PHQMD calculations are shown as colored bands.
@@ VI.3 Elliptic Flow
Now we step to a comparison of the PHQMD results for the elliptic flow for Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV with the STAR data [12]. In Fig. 21 we show the elliptic flow ($v_{2}$) of protons, deuterons and tritons, $^{3}\rm{He}$, and $^{4}\rm{He}$ as a function of the rapidity $y$ and integrated in the chosen $p_{T}$ ranges for the three EoS. For protons, $v_{2}$ is slightly negative in the considered $p_{T}$ interval. One can see that the $v_{2}$ of protons is quite well described by the SM EoS. The H EoS predicts almost zero $v_{2}$, while the S EoS predicts a positive $v_{2}$, which is opposite to the data. For deuterons, tritons, $^{3}\rm{He}$, and $^{4}\rm{He}$, the soft EoS yields a positive $v_{2}(y)$, in contradistinction to the negative $v_{2}$ of the experimental data. For all clusters SM and H EOS give rather similar results . Both show a more negative $v_{2}$ than the data in the whole rapidity range in distinction to the PHQMD calculations at $\sqrt{s_{\rm{NN}}}=2.4$ GeV [31], where the $v_{2}$ for SM and H is less negative than the data, however in a slightly different $p_{T}$ and centrality interval. It is remarkable that the sign of $v_{2}$ depends on the cluster size. For protons, $v_{2}$ is always negative, for $d$, $t$ and $^{3}{\rm{He}}$ it changes from positive values at y=0.5y=-0.5 to negative values close to midrapidity, whereas for heavier fragments $v_{2}$ is always positive. PHQMD calculations with H and SM reproduce the trend of $v_{2}$ as a function of the rapidity.
Figure 21: The elliptic flow $v_{2}$ of proton, deuteron, triton, $^{3}\rm{He}$, and $^{4}\rm{He}$ as a function of rapidity $10-40\%$ $\sqrt{s_{\rm{NN}}}=3$ GeV Au++Au collisions. The STAR data [14] are shown as colored markers while PHQMD calculations are shown as colored lines or bands.
@@
In the bottom right plot of Fig. 21 we compile $v_{2}(y)$ of the experimental results and the theoretical predictions with a SM EoS for protons, deuterons, triton, $^{3}\rm{He}$, and $^{4}\rm{He}$. The rapidity distributions of $v_{2}$ for different clusters exhibit a clear mass ordering, similar to $v_{1}$. This is qualitatively reproduced by the PHQMD calculations using a H or SM EoS. At mid-rapidity, 0.1<y<0-0.1<y<0, the value of $v_{2}$ is negative and nearly identical for $p$, $d$, and $^{3}\rm{He}$. A soft EoS give, as said, positive $v_{2}$ values for all clusters in 0.5<y<0-0.5<y<0, opposite to the data.
Figure 22: The elliptic flow $v_{2}$ of protons, deuterons, tritons, $^{3}\rm{He}$, and $^{4}\rm{He}$ in the rapidity regions (0.1,0)(-0.1,0) and (0.4,0.3)(-0.4,-0.3) as a function of $p_{T}$ in $10-40\%$ $\sqrt{s_{\rm{NN}}}=3$ GeV Au++Au collisions. The STAR data [14] are shown as black markers while PHQMD calculations are shown as colored bands.
The transverse momentum dependence of $v_{2}$ is presented in Fig. 22 for protons, deuterons, tritons, and $^{3}\rm{He}$ as a function of $p_{T}$ for the rapidity intervals 0.1<y<0-0.1<y<0 (left column) and 0.4<y<0.3-0.4<y<-0.3 (right column) for $10-40\%$ mid-central Au++Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. First of all we observe in the calculations a strong $p_{T}$ dependence of $v_{2}(p_{T})$ and a strong dependence on the EoS for all clusters and in both rapidity intervals. For both rapidity intervals the soft EoS gives mostly a positive $v_{2}$ for all clusters, whereas the data show mostly a negative $v_{2}$. Although the calculations using a H and a SM EoS are quite different for the proton $v_{2}(p_{T})$ , for deuterons, tritons and $^{3}\rm{He}$ the two EoS give almost the same result at forward as well as at mid-rapidity. Both give, however, a less positive or more negative $v_{2}$ compared to the experimental data.
For protons at midrapidity, the SM EoS gives the largest negative $v_{2}(p_{T})$, followed by the hard EoS, similar to what has been observed in $v_{2}(y)$. Calculations with SM give a result close to the experiment data up to $p_{T}=1.4$ GeV/c/c, but above 1.4 GeV/c/c, the calculations continue to decrease while the data increase. For all clusters H and SM calculations predict a similar $v_{2}(p_{T})$. The $v_{2}$ calculated by PHQMD follows qualitatively the experimental data but give a larger negative $v_{2}$.
For the backward rapidity interval (0.4<y<0.3-0.4<y<-0.3, Fig. 22, right column) the difference between theory and experiment is more pronounced. The proton data show a $v_{2}(p_{T})$, which is slightly less negative than that in the rapidity interval 0.1<y<0-0.1<y<0, whereas the PHQMD calculations yield a more negative $v_{2}$, similar to that observed at midrapidity. The cluster data display a qualitative different behaviour: The PHQMD results are quite similar to those obtained at midrapidity whereas the deuteron data show a $v_{2}(p_{T})$, which is compatible with zero, and larger clusters show a positive $v_{2}(p_{T})$. Thus the experimental $v_{2}$ of clusters changes sign when going from $\sqrt{s_{\rm{NN}}}=2.4$ to $\sqrt{s_{\rm{NN}}}=3$ GeV [28]. This indicates that in this rapidity interval clusters are more than an ensemble of randomly picked nucleons. The PHQMD calculations yield, as for $\sqrt{s_{\rm{NN}}}$ = 2.4 GeV, for $p_{T}>1$ GeV a negative $v_{2}(p_{T})$. This discrepancy between theory and experiment points towards a cluster production process, which is not yet correctly modeled in PHQMD and which will be explored in an upcoming publication.
Figure 23: The mass number scaled elliptic flow $v_{2}/A$ of protons, deuterons, tritons, and $^{3}\rm{He}$ as a function of $p_{T}/A$ in different rapidity intervals in $10-40\%$ mid-central Au++Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. The STAR data [14] are shown as colored markers while PHQMD SM EoS calculations are shown as colored bands.
Fig. 23 shows the compilation of elliptic flow, scaled by the mass number $v_{2}/A$, for protons, deuterons, tritons, and $^{3}\rm{He}$ in different rapidity intervals for the SM EoS at $\sqrt{s_{\rm{NN}}}=3$ GeV in comparison with the data. Data show although the $v_{2}$ of the light clusters approximately follows a mass number scaling at midrapidity, in data as well as in the calculations, the protons do not. At $\sqrt{s_{\rm{NN}}}=2.4$ [1] a scaling of $v_{2}/A(p_{2}/A)$ for protons and clusters has been observed. This scaling is predicted if one assumes that $A$ nucleons combine randomly to form a cluster of size $A$. The violation of this scaling indicates that at $\sqrt{s_{\rm{NN}}}$ = 3 GeV clusters are not randomly combined nucleons but that another cluster formation mechanism exist, which are only approximately captured in the PHQMD approach
However, in data the $v_{2}(p_{T})$ of protons is less than $v_{2}/A(p_{2}/A)$ of clusters, for the PHQMD calculations it is opposite. For the data the fact that $v_{2}/A(p_{2}/A)$ scales for clusters but is different from $v_{2}/A(p_{2}/A)$ for protons continues to larger negative rapidities whereas the PHQMD results of $v_{2}/A(p_{2}/A)$ for protons approaches with increasing negative rapidity that of the clusters.
As already seen in the last figure, the PHQMD calculations reproduce the experimental $v_{2}(p_{T})$ of protons up to $p_{T}\approx 1.4$ GeV/c/c but continue to decrease with increasing $p_{T}$ whereas in experiment $v_{2}(p_{T})$ of protons bends over.
In Fig. 24 we present the PHQMD results for $v_{2}(y)$ of protons and $\Lambda$ in comparison to the STAR data [14] for $\sqrt{s_{\rm{NN}}}=3$ GeV in $10-40\%$ central Au+Au collisions and for the $p_{T}$ interval $0.4<p_{T}<2.0$ GeV/c/c for calculations with a H, S and SM EoS. We see, first of all, that in experiment $v_{2}$ for $\Lambda$ and $p$ agree within error bars in the whole rapidity range. This is not the case for the PHQMD calculations. We observe that a soft EoS substantially overestimates the $v_{2}(y)$ for both, $\Lambda$ and proton, in this $p_{T}$ interval. The SM EoS gives a good description of the data. For H EoS, the proton $v_{2}(y)$ is overestimated, particularly at mid-rapidity, while for $\Lambda$ a H EoS can describe the data for $y\geq-0.6$, although this may be partly due to the uncertainty of the data. In PHQMD calculations, the proton $v_{2}$ is less negative than that of $\Lambda$ at mid-rapidity for all EoS, but the situation appears reversed near target rapidity. We note, however, that in the target rapidity domain proton emission from spectator remnants may occur in a time scale larger than the time for which we pursue the calculation.
Figure 24: The elliptic flow $v_{2}$ of proton and $\Lambda$ as a function of rapidity in $10-40\%$ $\sqrt{s_{\rm{NN}}}=3$ GeV Au++Au collisions. The STAR data [12] are shown as closed (proton) and open ($\Lambda$) markers while PHQMD calculations are shown as solid (proton) and dashed ($\Lambda$) lines.
VII Summary
In this study we have compared the results of PHQMD calculations with the most extensive data set (of 3 GeV Au+Au collisions) from the Beam Energy Scan Program phase II of the STAR collaboration, which is presently and in near future available in the energy range of the future FAIR accelerator. The goal was to see how transport calculations for hadrons and, more important, for clusters compare with these high statistics data. For this purpose we studied multiplicities, $p_{T}$ spectra, rapidity distributions as well as the in-plane ($v_{1}$) and elliptic ($v_{2}$) flow for baryons and clusters and applied three different equations of state, a static hard and soft EOS and a momentum dependent soft EoS.
We see, first of all, that at these energies different hadronic EoS gives quite different results for $v_{1}$ and $v_{2}$ but influence as well the $p_{T}$ distributions of elementary baryons and clusters.
Most of the $\sqrt{s_{\rm{NN}}}=3$ GeV Au+Au data are best described by calculations using a soft momentum dependent EoS (SM), which takes into account the measured momentum dependence of the optical potential.
[…]
The flow of hadrons and clusters shows some surprising results. The experimental in-plane and directed flow of $p$ and $\Lambda$ agrees within error bars whereas their radial flow differs. In PHQMD the difference in the radial flow is due to the fact that the production of strange hadrons happens more probable close to the center of the reaction. In PHQMD calculations this provokes also a difference in the in-plane flow.
The PHQMD calculations show that also at this energy the in-plane flow $v_{1}$ depends on the nuclear equation of state and it is assuring that different extrapolations of the optical potential from the measured beam energy region to the energies required for studying $\sqrt{s_{\rm{NN}}}=3$ GeV reactions do not influence the result for $v_{1}$. PHQMD calculations show the same scaling and the same functional form of $v_{1}/A(p_{T}/A)$ as seen in the data, but the numerical value of $v_{1}/A(p_{T}/A)$ is slightly smaller.
For the elliptic flow the data at $\sqrt{s_{\rm{NN}}}=3$ GeV do not show anymore the scaling in $v_{2}/A(p_{T}/A)$ observed at $\sqrt{s_{\rm{NN}}}=2.4$ GeV. This is a strong hint that at this energy the clusters are not anymore formed by a random selection of nucleons but that their production is a selective process. The scaling is still present for clusters but the protons deviate. The same is observed for PHQMD calculations. For $y\leq 0.2$ the data for $v_{2}/A(p_{T}/A)$ approaches zero for clusters whereas PHQMD calculations show still a finite negative $v_{2}$. The understanding of this observation will be the subject for a future investigation. One has also to take into account that different extrapolations of the optical potential to the energies of interest here influence $v_{2}$ in a non-negligible way.
The general qualitative and in many kinematical regions quantitative agreement between data and PHQMD calculations, even for the very sensitive flow observables, confirms the maturity of the transport approaches and that the underlying physical processes are to a large extent understood. The above mentioned discrepancies show as well where more intensive studies are necessary to improve these approaches as far as cluster and strangeness production is concerned.
APPENDIX A: